OpenStudy (anonymous):

Does this infinte series converge or diverge? Sigma from 1-infinity of 1n(n)/n^1.02

OpenStudy (anonymous):

n^(1.02)?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \ln(n)/n^1.02\]

OpenStudy (anonymous):

Use the integral test you will find it will converge.

OpenStudy (anonymous):

Are you familiar with integral test? Or should I go on?

OpenStudy (anonymous):

In case you aren't If \[\int\limits_{k}^{\infty} f(x) dx \] is convergent implies \[\sum_{k}^{\infty}\an] is convergent

OpenStudy (anonymous):

\[\sum_{k}^{\infty}an\] is convergent

OpenStudy (anonymous):

take \[f(x) = \frac{\ln(n)n^(1.02)}\] and k = 1

OpenStudy (anonymous):

f(x) = ln(n)/n^(1.02) take the integral from 1 to inf and show it does not diverge that is it has a finite answer. That will mean our series also converges. =]

OpenStudy (anonymous):

i am also interested in this problem, and i'd love to use the integral test, but i don't know how to integrate that. would you do it by parts?

OpenStudy (anonymous):

Yes do it by parts

OpenStudy (anonymous):

let u=ln(n), dv=n^(1.02)? see but from here what would v be?

OpenStudy (anonymous):

sorry, i don't mean to hijack the problem, it's just the same thing i've been having trouble with, too.

OpenStudy (anonymous):

oh and i meant dv=n^(-1.02)dx

OpenStudy (anonymous):

unless i'm setting u and dv to the wrong parts

OpenStudy (anonymous):

v is something stupid -50/n^(0.02)

OpenStudy (anonymous):

...what. where did the 50 come in?

OpenStudy (anonymous):

i can see the power rule making the exponent come down to 0.02, and also kicking the negative out front, but 50? where did that come from?

OpenStudy (anonymous):

1.02 = 1 + 1/50, sorry I hate decimals

OpenStudy (anonymous):

oh right, i see. thanks.

OpenStudy (anonymous):

Are we good?

OpenStudy (anonymous):

oh i am so good with this now. but then i definitely stole this question from the original poster, so... sorry about that, Aubun.

OpenStudy (anonymous):

Okay! I'm off to bed, have a big math test tomorrow

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Thank You